3 (AAR and GAR): The price of a security is $100 at time 0, $50 at time 1, and $100 at time 2. What’s the AAR and GAR?

4 (FETY): A tax-exempt bond has a yield of 6%, exempt from both federal taxes (30%) and state taxes (10%). What’s the FETY of a corporate bond on which both federal and state taxes apply?

5 (yields): A 90-day T-bills pays $10,000 at maturity, the current price is $9875. What’s the bank discount rate? The bond equivalent yield? Effective annual yield?

7. (CAPM): What’s the expected return for a security with beta=0.8? The market risk-premium is 4% and risk-free rate is 0.5%. Now the actual return of the security is 4%, is it a buy signal or a sell signal?
Answer #1
3) The periodic returns of the security are calculated below:
Time Value Return for the period
0 100
1 50 -50.00%
2 100 100.00%
AAR = (-50+100)/2 = 25%
GAR = ((1-0.5)*(1+1))^(1/2)-1 = 0.00%
4) FETY of a corporate bond = 0.06/(1-40%) = 10.00%
5) Bank discount rate = (10000-9875)/10000 = 1.25%
Discount rate per annum = 1.25%*90/360 = 5.00%
Bond equivalent yield = (125/9875)*365/90 = 5.13%
Effective annual yield = (1+125/9875)^(365/90)-1 = 5.23%
7) Expected return per CAPM = Risk free rate+Beta*Market risk premium
= 0.5%+0.8*4% = 3.70%
If the actual return of the security is 4%, it means that the
security is undervalued. It is a buy signal.